OpenStudy (anonymous):

sin (x+(pi/4))-sin (x-(pi/4))=1 Help to solve for solutions?

OpenStudy (anonymous):

one way would be to rewrite the left hand side as a single trig function it will take a raft of steps, but i think you get \[\sqrt{2}\cos(x)\]

OpenStudy (anonymous):

then it should be very easy you need to use the addition angle formula, and the subtraction angle formula for sine

OpenStudy (anonymous):

Sin (x+π/4)=sin x cos (π/4)+cos x sin (π/4) sin (x-π/4)=sin x cos (–π/4)-cos x sin( –π/4) Would using the sum/difference formulas help to solve? The question says they should be used but I'm not sure where to go from here.

OpenStudy (anonymous):

some of those are numbers right?

OpenStudy (anonymous):

\[\frac{\sqrt{2}}{2}\sin(x)+\frac{\sqrt{2}}{2}\cos(x)\] is the first line

OpenStudy (anonymous):

you made a mistake in the second line

OpenStudy (anonymous):

sin (x-π/4)=sin x cos (–π/4)-cos x sin( –π/4) should be sin (x-π/4)=sin x cos (π/4)-cos x sin( π/4)

OpenStudy (anonymous):

Yea, you're right I made a mistake with the - sign. How would I go about finding the solution?

OpenStudy (anonymous):

hmm now that i look more carefully, ;perhaps you get \[\sqrt{2}\sin(x)\] when you add

OpenStudy (anonymous):

no matter, now solve \[\sqrt{2}\sin(x)=1\] which is the same as \[\sin(x)=\frac{\sqrt2}{2}\]

OpenStudy (anonymous):

pi/4 and 7pi/4 ?